Write the point slope form of the equation of the line that passes through the points (5,-9) and (-6,1)
Answer by stanbon(75887)

Write the point slope form of the equation of the line that passes through the points (5,-9) and (-6,1)
---
slope = m = (1--9)/(-6-5) = 10/-11 = -10/11
-------------------
Form:: y = mx + b
Solve for "b"::
-9 = (-10/11)5 + b
-------
-9 + (50/11) = b
----
-99/11 + 50/11 = -49/11
-----
Equation:
y = (-10/11)x - (49/11)
-------------------
Cheers,
Stan H.
-------------